# The Nondecreasing Rank

### v1

@@ 5 Geometry of the Order Cone and the Existence of Nondecreasing Factorizations
When $\mathcal{C}_{j}=\mathcal{C}(\mathcal{P}_{j})$ in the above definition the tensor product of cones is the same as the set $\mathcal{N}_{<\infty}(\times_{i=1}^{k}\mathcal{P}_{j})$. The projective tensor product is a proper, polyhedral cone of dimension $\prod_{j=1}^{k}p_{j}$ whenever every $\mathcal{C}_{j}\subset\mathbb{R}^{p_{j}}$ is proper and polyhedral [40]. As the projective tensor product is equal to $\text{cone}(\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}:{\mathbf{v}}^{(j)}\in\mathcal{C}_{j})$ the extremal rays must be rank-one tensors. The following result [17, Thm 3.22] identifies all such extremal rays.
@@
If every $\mathcal{C}_{j}$ except for one is simplicial then $\big{(}\otimes_{j=1}^{k}\mathcal{C}_{j}\big{)}^{*}=\otimes_{j=1}^{k}\mathcal{C}_{j}^{*}$. If every $\mathcal{C}_{j}$ is also polyhedral then the facets of $\otimes_{j=1}^{k}\mathcal{C}_{j}$ have the form $\{{\mathbf{T}}:\langle{\mathbf{T}},\otimes_{j=1}^{k}{\mathbf{h}}^{(j)}\rangle=0\}\cap\otimes_{j=1}^{k}\mathcal{C}_{j}$ where every ${\mathbf{h}}^{(j)}$ is extremal in $\mathcal{C}_{j}^{*}$.
Using Lemma 1, a $\mathcal{V}$-representation of $\mathcal{N}_{<\infty}$ can be obtained from the extremal rays of each $\mathcal{C}(\mathcal{P}_{j})$. The next result [25, Thm 7] describes the extremal rays of any order cone in terms of upsets. An upset $\mathcal{U}\subseteq\mathcal{P}$ is a subset with the property that if $x\in\mathcal{U}$ and $x\preceq y$ then $y\in\mathcal{U}$. An upset is connected if the subgraph of the Hasse diagram corresponding to the upset is connected.
@@
Lemma 2.
Definition 4.
A collider in the Hasse diagram refers to a subgraph of the form
cab
Equivalently, a collider is a set of three elements $a,b,c$, where $a\lessdot c$ and $b\lessdot c$.
Theorem 2.
The cone $\mathcal{C}(\mathcal{P})$ is simplicial if and only if the Hasse diagram does not contain any colliders.
Proof.
If $\mathcal{P}$ has $p$ elements, the cone $\mathcal{C}(\mathcal{P})$ is simplicial if and only if the number of connected upsets of $\mathcal{P}$ is $p$. For every $x\in\mathcal{P}$, the intervals $x,\infty)=\{y:x\preceq y\}$ are distinct connected upsets. Thus, the number of connected upsets is at least $p$. To show that the order cone is simplicial it suffices to show that the number of connected upsets within each connected component of size $m$ in the Hasse diagram is equal to $m$.
Conversely, assume that $\mathcal{P}$ has a connected component of the Hasse diagram of size $m$ that has a collider. There exists three elements $a,b,c$ with $a,b$ incomparable and $a\prec c$, $b\prec c$. The upset $a\cup[c,\infty)$ is connected and does not equal $x,\infty)$ for any $x\in\mathcal{P}$. This shows that this connected component of $\mathcal{P}$ has at least $m+1$ connected upsets. ∎
Any equivalent phrasing of Theorem 2 is that the Hasse diagram is union of trees where within each tree there is a minimum element root node and every edge in the tree is directed away from the root node (See Figure 4). Figure 3 (a) and (b) respectively illustrate order cones that are and are not simplicial. In the order constrained statistical inference literature a few orderings that commonly occur are chains, the tree order where $x_{1}\prec x_{2},\ldots,x_{p}$, and the umbrella order where $x_{1}\prec\cdots\prec x_{l-1}\prec x_{l}\succ x_{l+1}\cdots\succ x_{p}$ [48, Sec 2.3]. Only the umbrella order out of these three orders is not simplicial.
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$$x_{7}$$x_{8}$
Figure 4: A Hasse diagram with a tree shape that has no colliders.
@@
When $\mathcal{P}=\times_{j=1}^{k}\mathcal{P}_{j}$ the cone $\mathcal{C}(\mathcal{P})$ is equal to $\mathcal{N}_{<\infty}$ if and only if every poset $\mathcal{P}_{j}$ except for one is trivial. A trivial poset has no ordering constraints apart from nonnegativity and has no arrows in its Hasse diagram. If every $\mathcal{P}_{j}=[p_{j}]$ is a chain, the cone $\mathcal{N}_{<\infty}$ is generated by $\prod_{j=1}^{k}p_{j}$ extremal rays, while $\mathcal{C}(\mathcal{P})$ has the same number of extremal rays as there are non-empty subsets $\mathcal{A}\subset\mathcal{P}$ where no two elements of $\mathcal{A}$ are comparable. In particular, when $k=2$ the number of extremal rays of $\mathcal{C}(\mathcal{P})$ is ${p_{1}+p_{2}\choose p_{1}}-1$.
@@
Let $\mathcal{P}_{j}$ have size $p_{j}$ and have $q_{j}$ connected upsets. The extremal rays of $\mathcal{N}_{<\infty}$ are all extremal rays of $\mathcal{C}(\mathcal{P})$ by Lemmas 1 and 2. Thus, $\mathcal{N}_{<\infty}=\mathcal{C}(\mathcal{P})$ if and only if these two cones have the same number of extremal rays. Assume that $\mathcal{P}_{1}$ is the only non-trivial poset. Then $\times_{j=1}^{k}\mathcal{P}_{j}$ is a disjoint union of $\prod_{j=2}^{k}p_{j}$ copies of the poset $\mathcal{P}_{1}$. The number of connected upsets of $\mathcal{P}$ is $q_{1}\prod_{j=2}^{k}p_{j}$ and is equal to the number of extremal rays of $\mathcal{C}(\mathcal{P})$. As $p_{j}=q_{j}$ for $j>1$, $q_{1}\prod_{j=2}^{k}p_{j}=\prod_{i=1}^{k}q_{j}$ is equal to the product of the number of connected upsets of each $\mathcal{P}_{j}$, namely the number of extremal rays of $\mathcal{N}_{<\infty}$.
To show the other direction, assume that the inequalities $a_{j}\prec b_{j}$ are present in $\mathcal{P}_{j}$ for $j=1,2$. If $\mathcal{U}_{j}$ is a connected upset in $\mathcal{P}_{j}$ then $\times_{j=1}^{k}\mathcal{U}_{j}$ is a connected upset in $\mathcal{P}$. Letting $x_{j}$ be a maximal element of $\mathcal{P}_{j}$, $j=3,\ldots,k$, the set
@@
To prove the last result we count the number of connected upsets of $\times_{j=1}^{k}[p_{j}]$. This is equal to the number of non-empty antichains — sets where no two elements in a set are comparable. There is a bijection between antichains $\mathcal{A}\subset\mathcal{P}$ and connected upsets $\cup_{a\in\mathcal{A}}[a,\infty)$. If $\mathcal{A}=\{a_{1},\ldots,a_{m}\}$ are the minimal elements of $\mathcal{U}$ then $\mathcal{A}$ is an antichain with $\mathcal{U}=\cup_{a\in\mathcal{A}}[a,\infty)$ and conversely any antichain $\mathcal{A}$ produces the upset $\cup_{a\in\mathcal{A}}[a,\infty)$ with minimal elements $\mathcal{A}$. This upset, and in fact any upset, is connected when the poset contains a maximum element, as is the case when $\mathcal{P}=\times_{j=1}^{k}[p_{j}]$ is a product of chains. When $k=2$ the number of antichains is equal to the number of staircase matrices. A staircase matrix is determined by a path with non-decreasing coordinates in the lattice $\mathbb{N}\times\mathbb{N}$ from $(0,0)$ to $(p_{1},p_{2})$. Such a path must have length $p_{1}+p_{2}$, and out of the $p_{1}+p_{2}$ steps along the path, $p_{1}$ of these steps must be upward in the direction $(1,0)$. The number of such paths is $p_{1}+p_{2}\choose p_{1}$}. As the path $(0,0)-(1,0)-\cdots-(p_{1},0)-(p_{1},1)-\cdots-(p_{1},p_{2})$ corresponds to the $\mathbf{0}$} staircase matrix, or the empty antichain, we must remove this path from consideration. The number of staircase matrices is also known as the number of Dyck paths from $(0,0)$ to $(p_{1},p_{2})$. ∎
@@
An implication of Theorem 3 is that if a random tensor $\mathbf{T}$} is drawn from a measure that is absolutely continuous with respect to the Lebesgue measure on $\mathcal{C}(\mathcal{P})$ then $\text{Pr}({\mathbf{T}}\in\mathcal{N}_{<\infty})>0$ and $\text{Pr}({\mathbf{T}}\in\mathcal{C}(\mathcal{P})\backslash\mathcal{N}_{<\infty})>0$. The probability that a $\mathbf{T}$} drawn from the uniform distribution on the order polytope $\mathcal{O}(\mathcal{P})$ lies in $\mathcal{N}_{<\infty}\cap\mathcal{O}(\mathcal{P})$ when $\mathcal{P}=[m]\times[m]$ is found to equal $50\%,2.38\%$, and $0.004\%$ when $m=2,3,4$ by a volume computation. As the dimension of the matrix grows, the number of extremal rays of $\mathcal{C}(\mathcal{P})$ becomes significantly larger than that of $\mathcal{N}_{<\infty}$, and the probability of observing a random monotone tensor with finite ND rank is small. This suggests that when a data matrix ${\mathbf{T}}={\mathbf{X}}+\bm{\epsilon}$ is observed by taking a finite ND rank matrix $\mathbf{X}$} and perturbing it by noise $\bm{\epsilon}$, there is only a small probability that the observed $\mathbf{T}$} has finite ND rank. From an parameter estimation perspective this does not pose a serious issue as the observed $\mathbf{T}$} can be projected onto $\mathcal{N}_{<\infty}$ or $\mathcal{N}_{\leq r}$. We consider such low-rank approximations in Section 10
@@
Given a tensor $\mathbf{T}$}, it must first be ascertained that ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ if $\mathbf{T}$} is to have an exact ND factorization. Equations for determining if ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ can be found algorithmically by converting the $\mathcal{V}$-representation of $\mathcal{N}_{<\infty}$ in Lemma 1 into a $\mathcal{H}$-representation. In the next section we provide simple equations for testing membership when every $\mathcal{P}_{i}$ except for one has a tree structure. Below, the $24$ equations of $\mathcal{N}_{<\infty}\subset\mathbb{R}_{+}^{3\times 3}$ are computed in Polymake [26] for the case where $\mathcal{P}_{i}=\{a^{(i)},b^{(i)},c^{(i)}\}$, $i=1,2$ is given the collider structure specified in Definition 4, and the third row and column respectively correspond to the maximum elements $c^{(1)},c^{(2)}$.
@@
The first three rows of equations correspond to simple positivity and monotonicity constraints. The next two rows of equations have four non-zero variables and will appear in the next section as the type of equations required to cut-out $\mathcal{N}_{<\infty}$ when the posets are trees. The final four rows of equations are more exotic as they have dual vectors that are matrices with rank greater than one. For example, the last two inequalities have the corresponding dual vectors
@@
These dual vectors do not lie in $\otimes_{j=1}^{2}\mathcal{C}(\mathcal{P}_{j})^{*}$, for if they did they would be extremal in $\otimes_{j=1}^{2}\mathcal{C}(\mathcal{P}_{j})^{*}$ as they are normal vectors of facets of $\otimes_{j=1}^{2}\mathcal{C}(\mathcal{P}_{j})$, and by Lemma 1 they would be rank one matrices.

### v2

@@ 5 Geometry of the Order Cone and the Existence of Nondecreasing Factorizations
When $\mathcal{C}_{j}=\mathcal{C}(\mathcal{P}_{j})$ in the above definition the tensor product of cones is the same as the set $\mathcal{N}_{<\infty}(\times_{i=1}^{k}\mathcal{P}_{j})$. The projective tensor product is a proper, polyhedral cone of dimension $\prod_{j=1}^{k}p_{j}$ whenever every $\mathcal{C}_{j}\subset\mathbb{R}^{p_{j}}$ is proper and polyhedral [41]. As the projective tensor product is equal to $\text{cone}(\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}:{\mathbf{v}}^{(j)}\in\mathcal{C}_{j})$ the extremal rays must be rank-one tensors. The following result [18, Thm 3.22] identifies all such extremal rays.
@@
If at most one of the $\mathcal{C}_{j}$ cones is not simplicial then $\big(\otimes_{j=1}^{k}\mathcal{C}_{j}\big)^{*}=\otimes_{j=1}^{k}\mathcal{C}_{j}^{*}$. If every $\mathcal{C}_{j}$ is also polyhedral then the facets of $\otimes_{j=1}^{k}\mathcal{C}_{j}$ have the form $\{{\mathbf{T}}:\langle{\mathbf{T}},\otimes_{j=1}^{k}{\mathbf{h}}^{(j)}\rangle=0\}\cap\otimes_{j=1}^{k}\mathcal{C}_{j}$ where every ${\mathbf{h}}^{(j)}$ is extremal in $\mathcal{C}_{j}^{*}$.
Using Lemma 1, a $\mathcal{V}$-representation of $\mathcal{N}_{<\infty}$ can be obtained from the extremal rays of each $\mathcal{C}(\mathcal{P}_{j})$. The next result [26, Thm 7] describes the extremal rays of any order cone in terms of upsets. An upset $\mathcal{U}\subseteq\mathcal{P}$ is a subset with the property that if $x\in\mathcal{U}$ and $x\preceq y$ then $y\in\mathcal{U}$. An upset is connected if the subgraph of the Hasse diagram corresponding to the upset is connected.
Lemma 2 ([26]).
Definition 4.
A collider in the Hasse diagram of a poset $\mathcal{P}$ refers to a subgraph of the form
cab
Equivalently, a collider is a set of three elements $a,b,c\in\mathcal{P}$ where $a\lessdot c$ and $b\lessdot c$.
Theorem 2.
The cone $\mathcal{C}(\mathcal{P})$ is simplicial if and only if the Hasse diagram does not contain any colliders. A Hasse diagram has no colliders if and only if as an undirected graph it is a disjoint union of trees where each tree has edges directed away from one its nodes.
Proof.
If $\mathcal{P}$ has $p$ elements, by Lemma 2, the cone $\mathcal{C}(\mathcal{P})$ is simplicial if and only if the number of connected upsets of $\mathcal{P}$ is $p$. For every $x\in\mathcal{P}$, the intervals $x,\infty)=\{y:x\preceq y\}$ are distinct connected upsets. Thus, the number of connected upsets is at least $p$. To show that the order cone is simplicial it suffices to show that the number of connected upsets within each connected component of size $m$ in the Hasse diagram is equal to $m$.
Conversely, assume that $\mathcal{P}$ has a connected component of the Hasse diagram of size $m$ that has a collider. There exists three elements $a,b,c$ with $a,b$ incomparable and $a\prec c$, $b\prec c$. The upset $a\cup b\cup[c,\infty)$ is connected and does not equal $x,\infty)$ for any $x\in\mathcal{P}$. This shows that this connected component of $\mathcal{P}$ has at least $m+1$ connected upsets.
The last statement that the absence of colliders implies the given tree structure is proven by noting that if the undirected graph associated with the Hasse diagram contains a cycle, the only way to orient the edges in a cycle to avoid the appearance of a collider is to create a directed cycle. This contradicts the fact that Hasse diagrams are directed acyclic graphs. If a tree does not contain a minimum element that all edges are directed away from there exists at least two distinct minimal elements $a,b$ in the tree. As the tree is connected there exists a path between $a$ and $b$ with arrows directed away from both $a$ and $b$. There is no way to orient the edges of such a path without obtaining a collider. The converse direction that unions of trees with the given structure contain no colliders is apparent. ∎
Figure 3 illustrates the tree structure described in Theorem 2, while Figure 4 (a) and (b) respectively illustrate order cones that are and are not simplicial. In the order constrained statistical inference literature a few orderings that commonly occur are chains, the tree order where $x_{1}\prec x_{2},\ldots,x_{p}$, and the umbrella order where $x_{1}\prec\cdots\prec x_{l-1}\prec x_{l}\succ x_{l+1}\cdots\succ x_{p}$ [49, Sec 2.3]. The umbrella order represents a kind of unimodality where the mode is known to occur at $x_{l}$. Only the umbrella order out of these three orders is not simplicial.
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}$$x_{7}$$x_{8}$
Figure 3: A Hasse diagram with a tree shape that has no colliders.
@@
When $\mathcal{P}=\times_{j=1}^{k}\mathcal{P}_{j}$ the cone $\mathcal{C}(\mathcal{P})$ is equal to $\mathcal{N}_{<\infty}$ if and only if every poset $\mathcal{P}_{j}$ except for one is trivial. A trivial poset has no arrows in its Hasse diagram. If every $\mathcal{P}_{j}=[p_{j}]=\{1,\ldots,p_{j}\}$ is a chain, the cone $\mathcal{N}_{<\infty}$ is generated by $\prod_{j=1}^{k}p_{j}$ extremal rays, while $\mathcal{C}(\mathcal{P})$ has the same number of extremal rays as there are non-empty subsets $\mathcal{A}\subset\mathcal{P}$, where no two elements of $\mathcal{A}$ are comparable. In particular, when $k=2$ the number of extremal rays of $\mathcal{C}(\mathcal{P})$ is ${p_{1}+p_{2}\choose p_{1}}-1$.
@@
Let $\mathcal{P}_{j}$ have size $p_{j}$ and have $q_{j}$ connected upsets. The extremal rays of $\mathcal{N}_{<\infty}$ are all extremal rays of $\mathcal{C}(\mathcal{P})$. To see this, observe that every extremal ray of $\mathcal{N}_{<\infty}$ has the form $\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$ where ${\mathbf{v}}^{(j)}$ is extremal in $\mathcal{C}(\mathcal{P}_{j})$ by Lemma 1. By Lemma 2 each ${\mathbf{v}}^{(j)}$ is associated with a connected upset $\mathcal{U}_{j}\subset\mathcal{P}_{j}$. Any product $\times_{j=1}^{k}\mathcal{U}_{j}$ of connected upsets is a connected upset in $\mathcal{P}$. As $\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$ is associated with the connected upset $\times_{j=1}^{k}\mathcal{U}_{j}$, another invocation of Lemma 2 shows that $\otimes_{j=1}^{k}{\mathbf{v}}^{(j)}$ is extremal in $\mathcal{C}(\mathcal{P})$.
As a consequence, $\mathcal{N}_{<\infty}=\mathcal{C}(\mathcal{P})$ if and only if these two cones have the same number of extremal rays. Assume that $\mathcal{P}_{1}$ is the only non-trivial poset. Then $\times_{j=1}^{k}\mathcal{P}_{j}$ is a disjoint union of $\prod_{j=2}^{k}p_{j}$ copies of the poset $\mathcal{P}_{1}$. The number of connected upsets of $\mathcal{P}$ is $q_{1}\prod_{j=2}^{k}p_{j}$ and is equal to the number of extremal rays of $\mathcal{C}(\mathcal{P})$. As $p_{j}=q_{j}$ for $j>1$, $q_{1}\prod_{j=2}^{k}p_{j}=\prod_{i=1}^{k}q_{j}$ is equal to the product of the number of connected upsets of each $\mathcal{P}_{j}$, namely the number of extremal rays of $\mathcal{N}_{<\infty}$.
To show the other direction, assume that the inequalities $a_{j}\prec b_{j}$ are present in $\mathcal{P}_{j}$ for $j=1,2$. Letting $x_{j}$ be a maximal element of $\mathcal{P}_{j}$, $j=3,\ldots,k$, the set
@@
To prove the last result we count the number of connected upsets of $\times_{j=1}^{k}[p_{j}]$. This is equal to the number of non-empty antichains — sets where no two elements in a set are comparable. There is a bijection between antichains $\mathcal{A}\subset\mathcal{P}$ and connected upsets $\cup_{a\in\mathcal{A}}[a,\infty)$. If $\mathcal{A}=\{a_{1},\ldots,a_{m}\}$ are the minimal elements of $\mathcal{U}$ then $\mathcal{A}$ is an antichain with $\mathcal{U}=\cup_{a\in\mathcal{A}}[a,\infty)$ and conversely any antichain $\mathcal{A}$ produces the upset $\cup_{a\in\mathcal{A}}[a,\infty)$ with minimal elements $\mathcal{A}$. This upset, and in fact any upset, is connected when the poset contains a maximum element, as is the case when $\mathcal{P}=\times_{j=1}^{k}[p_{j}]$ is a product of chains. When $k=2$ the number of antichains is equal to the number of staircase matrices; if $(x_{1},y_{1}),\ldots,(x_{l},y_{l})$ is an antichain then this determines a unique staircase matrix $\mathbf{M}$}, where $M_{ab}=1$ if and only if there exists an $i=1,\ldots,l$ with $(a,b)\succeq(x_{i},y_{i})$. A staircase matrix is also determined by a path with non-decreasing coordinates in the lattice $\mathbb{N}\times\mathbb{N}$ from $(0,0)$ to $(p_{1},p_{2})$. Such a path must have length $p_{1}+p_{2}$, and out of the $p_{1}+p_{2}$ steps along the path, $p_{1}$ of these steps must be upward in the direction $(1,0)$. The number of such paths is $p_{1}+p_{2}\choose p_{1}$}. As the path $(0,0)-(1,0)-\cdots-(p_{1},0)-(p_{1},1)-\cdots-(p_{1},p_{2})$ corresponds to the $\mathbf{0}$} staircase matrix, or the empty antichain, we must remove this path from consideration. ∎
@@
An implication of Theorem 3 is that if a random tensor $\mathbf{T}$} is drawn from a measure that is absolutely continuous with respect to the Lebesgue measure on $\mathcal{C}(\mathcal{P})$ then $\text{Pr}({\mathbf{T}}\in\mathcal{N}_{<\infty})>0$ and $\text{Pr}({\mathbf{T}}\in\mathcal{C}(\mathcal{P})\backslash\mathcal{N}_{<\infty})>0$. The probability that a $\mathbf{T}$} drawn from the uniform distribution on the order polytope $\mathcal{O}(\mathcal{P})$ lies in $\mathcal{N}_{<\infty}\cap\mathcal{O}(\mathcal{P})$ when $\mathcal{P}=[m]\times[m]$ is found to equal $50\%,2.38\%$, and $0.004\%$ when $m=2,3,4$ by a volume computation. As the dimension of the matrix grows, the number of extremal rays of $\mathcal{C}(\mathcal{P})$ becomes significantly larger than that of $\mathcal{N}_{<\infty}$, and the probability of observing a random monotone tensor with finite ND rank is small. This suggests that when a data matrix ${\mathbf{T}}={\mathbf{X}}+\bm{\epsilon}$ is observed by taking a finite ND rank matrix $\mathbf{X}$} and perturbing it by noise $\bm{\epsilon}$, there is only a small probability that the observed $\mathbf{T}$} has finite ND rank. From a parameter estimation perspective this does not pose a serious issue as the observed $\mathbf{T}$} can be projected onto $\mathcal{N}_{<\infty}$ or $\mathcal{N}_{\leq r}$. We consider such low-rank approximations in Section 10.
@@
Given a tensor $\mathbf{T}$}, it must first be ascertained that ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ if $\mathbf{T}$} is to have an exact ND factorization. Equations for determining if ${\mathbf{T}}\in\mathcal{N}_{<\infty}$ can be found algorithmically by converting the $\mathcal{V}$-representation of $\mathcal{N}_{<\infty}$ in Lemma 1 into a $\mathcal{H}$-representation. In the next section we provide simple equations for testing membership when every $\mathcal{P}_{i}$ except for one has a tree structure. Below, the $24$ equations of $\mathcal{N}_{<\infty}\subset\mathbb{R}_{+}^{3\times 3}$ are computed in Polymake [27] for the case where $\mathcal{P}_{i}=\{a^{(i)},b^{(i)},c^{(i)}\}$, $i=1,2$ is given by the collider structure specified in Definition 4, and the third row and column respectively correspond to the maximum elements $c^{(1)},c^{(2)}$.
@@
The first three rows of equations correspond to simple positivity and monotonicity constraints. The next two rows of equations have four non-zero variables and will appear in the next section as the type of equations required to cut-out $\mathcal{N}_{<\infty}$ when the posets are trees. The final four rows of equations are more exotic as they can be written as $\text{tr}({\mathbf{C}}^{\intercal}{\mathbf{T}})\leq 0$ where the matrix $\mathbf{C}$} has a rank greater than one. For example, the last two inequalities have the corresponding $\mathbf{C}$} matrices
@@
These matrices do not lie in $\otimes_{j=1}^{2}\mathcal{C}(\mathcal{P}_{j})^{*}$, for if they did they would be extremal in $\otimes_{j=1}^{2}\mathcal{C}(\mathcal{P}_{j})^{*}$ as they are normal vectors of facets of $\otimes_{j=1}^{2}\mathcal{C}(\mathcal{P}_{j})$, and by Lemma 1 they would be rank one matrices.